# Assessing the Macro and Micro Effects of Random Seeds on Fine-Tuning Large Language Models

### v1

@@ 6 Conclusion
In conclusion, this work highlights the significant impact of random seeds on pretrained LLMs, revealing variability at both macro and micro levels. By introducing a novel consistency metric, we emphasize the importance of considering seed-induced variations in individual predictions in model evaluation. Our findings stress the need for incorporating random seed sensitivity into benchmarking for more reliable and reproducible results.
@@ 7 Limitations
Due to limited computing resources, our experiments were conducted with only five random seeds, which may not be sufficient for drawing broader generalizations of the findings and implications. Additionally, the reference SOTA scores for GLUE and SuperGLUE tasks were obtained using the pretrained LLM RoBERTa-large, and therefore, we conducted experiments solely on RoBERTa-large. Expanding the experiments to include various LLMs, particularly larger-scale models, would strengthen our findings and conclusions. Furthermore, incorporating more NLP benchmark datasets would provide a more comprehensive evaluation, as diverse datasets would better capture variability across tasks, domains, and data distributions, ultimately enhancing the robustness and applicability of our analysis. Additionally, our findings and implications are more suited for classification tasks, as only 1 out of the 13 tasks in our experiments is a regression task. Therefore, more comprehensive experiments should be conducted specifically on various regression tasks. The ACC-CON metric, which is not directly applicable to regression tasks, hinders the ability to evaluate correct consistency in this context.
@@ Appendix A Appendix
A.1 Data Description
@@
A.2 Hyperparameter Settings
Table 4 provides the detailed hyperparameter configurations. Unless stated otherwise, we adopt the default hyperparameter values from the Hugging Face framework.
A.3 Replicated SOTA Scores
To ensure the reproducibility of our experiments in SuperGLUE and GLUE tasks, we adhered to the specified settings and reproduced the state-of-the-art (SOTA) accuracy scores reported in: https://github.com/facebookresearch/fairseq/tree/main/examples/roberta. Our replicated accuracy scores for the GLUE and SuperGLUE tasks, presented in Table 2, are directly comparable and align with those shown in Table 1 of the main paper and Table 5 in Section A.4.
A.4 Additional Results
Table 5 presents model performance across various metrics, including precision (P), recall (R), F1 score, accuracy, CON, and ACC-CON, with average values and standard deviations (VAR). In Section 5 of the main paper, significant variance in macro-level performance across many tasks highlights sensitivity to random seed selection. Similar patterns in the VAR values for P, R, and F1 further confirm the robustness of our findings across various standard metrics.
GLUE MRPC QNLI QQP SST2 RTEG COLA STSB
Reference 90.9 94.7 92.2 96.4 86.6 68.0 92.4
Replicated 91.2 94.7 92.1 96.9 84.8 65.3 92.5
@@ A.4 Additional Results
GLUE MRPC QNLI QQP SST2 RTEG COLA STSB
Classes 2 2 2 2 2 2 -
Train samples 3668 104743 363846 67349 2490 8551 5749
Dev samples 408 5463 40430 872 277 1043 1500
Test samples 1725 5463 39096 1821 3000 1063 1379
@@
Train samples 9427 250 2500 27243 5428 400
Dev samples 1886 50 500 4848 1200 100
Test samples 3270 57 278 953 638 500
@@
GLUE MRPC QNLI QQP SST2 RTEG COLA STSB
Random seed 42 72 42 52 52 72 42
Batch size 10 10 10 10 10 10 32
Epoch 8 6 8 7 10 8 3
Learning rate 2e-5 2e-5 1e-5 2e-5 1e-5 1e-5 4e-5
Learning rate schedule type linear linear linear linear linear linear linear
Max sequence length 512 512 512 512 512 512 512
Gradient accumulation steps 2 2 2 2 2 2 2
@@
Table 4: The hyperparameter settings for GLUE and SuperGLUE tasks to replicate the reference performance in Table 2.
@@
Tasks P R F1 Accuracy CON ACC-CON Tasks P R F1 Accuracy CON ACC-CON
MRPC 91.67 94.48 93.04 90.34 92.21 52.33 BoolQ 87.69 88.36 88.03 85.05 94.78 82.44
($\pm$0.47) ($\pm$1.9) ($\pm$0.75) ($\pm$0.93) ($\pm$0.98) ($\pm$1.19) ($\pm$0.39) ($\pm$0.32) ($\pm$0.18) ($\pm$0.24) ($\pm$0.28) ($\pm$0.24)
QNLI 95.47 93.62 94.53 94.53 96.84 92.95 CB 99.13 98.26 98.67 98.8 91.61 90.18
($\pm$0.28) ($\pm$0.31) ($\pm$0.17) ($\pm$0.16) ($\pm$0.3) ($\pm$0.21) ($\pm$1.95) ($\pm$2.38) ($\pm$1.22) ($\pm$1.1) ($\pm$2.53) ($\pm$2.95)
QQP 87.67 91.17 89.38 92.02 96.00 90.03 RTES 76.31 65.34 76.34 69.6 71.84 56.61
($\pm$0.36) ($\pm$0.53) ($\pm$0.11) ($\pm$0.07) ($\pm$0.07) ($\pm$0.04) ($\pm$19.45) ($\pm$38.38) ($\pm$9.12) ($\pm$18.22) ($\pm$17.63) ($\pm$14.4)
SST2 95.56 95.68 95.6 95.73 98.1 94.58 MultiRC 79.81 68.24 82.46 79.01 76.16 67.09
($\pm$0.57) ($\pm$0.89) ($\pm$0.36) ($\pm$0.71) ($\pm$0.27) ($\pm$0.27) ($\pm$0.66) ($\pm$38.17) ($\pm$0.65) ($\pm$12.21) ($\pm$18.87) ($\pm$16.48)
RTEG 87.57 77.25 82.08 84.04 92.78 80.43 WiC 65.12 79.06 71.28 68.4 79.09 57.95
($\pm$1.28) ($\pm$1.13) ($\pm$0.52) ($\pm$0.47) ($\pm$1.09) ($\pm$0.74) ($\pm$1.29) ($\pm$8.29) ($\pm$4.17) ($\pm$2.83) ($\pm$7.00) ($\pm$4.99)
COLA - - - 64.51 94.23 82.52 COPA 70.20 72.20 71.00 73.20 67.60 57.00
- - - ($\pm$0.64) ($\pm$0.64) ($\pm$0.35) ($\pm$14.46) ($\pm$13.79) ($\pm$13.55) ($\pm$12.83) ($\pm$10.29) ($\pm$11.98)
STSB - - - 92.19 98.53 -
- - - ($\pm$0.30) ($\pm$0.16) -
Table 5: Evaluation metrics used in this study. Accuracy is employed for all tasks except STSB and CoLA, where Pearson correlation and Matthew’s correlation coefficient are used, respectively. CON - consistency, ACC-CON - correct consistency.

### v2

@@ 6 Conclusion
In conclusion, this work highlights the significant impact of random seeds on pretrained LLMs, revealing variability at both macro and micro levels. By introducing a novel consistency metric, we emphasize the importance of considering seed-induced variability in individual predictions in model evaluation. Our findings stress the need for incorporating random seed sensitivity into benchmarking for more reliable and reproducible results.
@@ 7 Limitations
Our work focuses on classification tasks which are the main consists of GLUE and SuperGLUE benchmarks. While this provides a solid foundation, it may limit the generalizability of our findings to other NLP task types. Incorporating a broader set of benchmark datasets would allow for a more comprehensive evaluation using our proposed macro and micro metrics across diverse task categories (as summarized in Table 2). Experimenting with greater task diversity would better capture variability in model behavior, ultimately enhancing the robustness and applicability of our analysis. We leave this broader exploration as future work.
Acknowledgement
The authors would like to thank the anonymous reviewers for feedback that improved the paper, the New Jersey Institute of Technology (NJIT) and the US National Institutes of Health (NIH) for providing funding. This research is supported by NJIT New Faculty Grant and NIH Grant R01LM013486. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of NJIT or NIH.
@@ Appendix A Appendix
A.1 Macro and Micro Metrics for Various NLP Tasks
Given two LLMs $A$ and $B$, a dataset $\mathcal{D}={d_{1},\cdots,d_{N}}$ of $N$ data points, $y_{i}^{A}$ and $y_{i}^{B}$ are the prediction of AA,$B$ for a data point $d_{i}$ with ground truth $r_{i}$. Table 2 summarizes the performance metric $\zeta$ and the corresponding scoring functions $\pi(\cdot)$ used in Equations 2 and 3 for various NLP tasks.
A.2 Data Description
@@
A.3 Hyperparameter Settings
Table 5 and Table 6 provides the detailed hyperparameter configurations. Unless stated otherwise, we adopt the default hyperparameter values from the Hugging Face framework.
A.4 Replicated SOTA Scores
To ensure the reproducibility of our experiments in SuperGLUE and GLUE tasks, we adhered to the specified settings and reproduced the state-of-the-art (SOTA) accuracy scores reported in: https://github.com/facebookresearch/fairseq/tree/main/examples/roberta. Although differences in the fine-tuning script and missing settings from the original authors prevented us from reproducing the exact SOTA scores, our replicated accuracy scores for the GLUE and SuperGLUE tasks, presented in Table 3, are directly comparable and align with those shown in Table 1 of the main paper and Table 7 in Section A.5.
We were unable to replicate the representation (special token extractions) and model settings (unpublished pretrained model) for the WSC and MNLI tasks, so they are omitted from the experiment. STSB is a regression task, thus is omitted.
A.5 Additional Results
Table 7 and Table 8 present model performance across various metrics, including precision (P), recall (R), F1 score, accuracy, CON, and CCON, with average values and standard deviations (VAR). In Section 5 of the main paper, significant variance in macro-level performance across many tasks highlights sensitivity to random seed selection. Similar patterns in the VAR values for P, R, and F1 further confirm the robustness of our findings across various standard metrics.
A.6 Additional Discussion
Is there a universally superior random seed?
To investigate whether a specific random seed consistently leads to better results across different models or tasks, in Figure 3 we present a heatmap of normalized accuracy for each task across ten random seeds using RoBERTa-large. There is no significant difference in color distribution between each row, indicating that no discernible pattern or evidence supporting the existence of a universally superior random seed.
Figure 3: A heatmap of normalized $\zeta$ values across tasks and ten random seeds, with a darker color representing a better accuracy.
Will increasing training data size improve variance and consistency in general?
Figure 4 presents the correlation analysis of VAR, CON, and CCON using Llama3.2-3B, revealing patterns consistent with the findings discussed in Section 5.3.
Figure 4: Correlation between training size (log scale), VAR, CON, and CCON while using Llama3.2-3B. Tasks are arranged in ascending order of training size, with exact sizes detailed in Appendix 4.
NLP Task Type Standard Metric $\zeta$ Scoring Function $\pi(y_{i}^{A},y_{i}^{B})$ Scoring Function $\pi(y_{i}^{A},y_{i}^{B},r_{i})$
Classification Accuracy, F1 $\mathbf{1}(y_{i}^{A}=y_{i}^{B})$ $\mathbf{1}(y_{i}^{A}=y_{i}^{B}=r_{i})$
Regression Pearson/Spearman, MSE, MAE $\zeta(y^{A},y^{B})$ $\frac{1}{2}(\zeta(y_{i}^{A},r_{i})+\zeta(y_{i}^{B},r_{i}))$
Text Generation BLEU, ROUGE, BERTScore $\zeta(y_{i}^{A},y_{i}^{B})$ $\frac{1}{2}(\zeta(y_{i}^{A},r_{i})+\zeta(y_{i}^{B},r_{i}))$
Sequence Labeling Token-/Span-level F1 $\frac{1}{T}\sum_{t}\mathbf{1}(y_{i,t}^{A}=y_{i,t}^{B})$ $\frac{1}{T}\sum_{t}\mathbf{1}(y_{i,t}^{A}=y_{i,t}^{B}=r_{i,t})$
Structured Prediction UAS, LAS, Exact Match $\zeta(y_{i}^{A},y_{i}^{B})$ $\frac{1}{2}(\zeta(y_{i}^{A},r_{i})+\zeta(y_{i}^{B},r_{i}))$
Question Answering Exact Match, F1 (token overlap) $\zeta(y_{i}^{A},y_{i}^{B})$ $\frac{1}{2}(\zeta(y_{i}^{A},r_{i})+\zeta(y_{i}^{B},r_{i}))$
Retrieval / Ranking NDCG, MRR, MAP $\zeta(y_{i}^{A},y_{i}^{B})$ $\frac{1}{2}(\zeta(y_{i}^{A},r_{i})+\zeta(y_{i}^{B},r_{i}))$
Table 2: Common evaluation metrics and scoring functions for various NLP task types. $\mathbf{1}(\cdot)$ denotes the indicator function that equals 1 or 0. For generation and ranking tasks, similarity and ranking correlation metrics (like BERTScore or Kendall Tau) are more appropriate. Note that this table includes only a subset of commonly used $\zeta$ metrics; many other relevant metrics are omitted due to space limitations.
GLUE MRPC QNLI QQP SST2 RTEG CoLA
Reference 90.9 94.7 92.2 96.4 86.6 68.0
Replicated 91.2 94.7 92.1 96.9 84.8 65.3
@@ A.6 Additional Discussion
GLUE MRPC QNLI QQP SST2 RTEG CoLA
Classes 2 2 2 2 2 2
Train samples 3668 104743 363846 67349 2490 8551
Dev samples 408 5463 40430 872 277 1043
Test samples 1725 5463 390965 1821 3000 1063
@@
Train samples 9427 250 2490 27243 5428 400
Dev samples 3270 56 277 4848 638 100
Test samples 3245 250 3000 9693 1400 500
@@
GLUE MRPC QNLI QQP SST2 RTEG CoLA
Random seed 42 72 42 52 52 72
Batch size 10 10 10 10 10 10
Epoch 8 6 8 7 10 8
Learning rate 2e-5 2e-5 1e-5 2e-5 1e-5 1e-5
Max sequence length 512 512 512 512 512 512
Gradient accumulation steps 2 2 2 2 2 2
@@
Table 5: The hyperparameter settings for GLUE and SuperGLUE tasks to replicate the reference performance in Table 3 using RoBERTa-large.
GLUE MRPC QNLI QQP SST2 RTEG CoLA
Batch size 10 10 32 10 10 10
Epoch 20 6 2 6 20 6
Learning rate 3e-5 3e-5 2e-5 3e-5 3e-5 1e-5
Max sequence length 512 512 512 512 512 512
Gradient accumulation steps 2 2 2 2 2 2
LoRA $r_{q},r_{v}$ 8 8 8 8 8 8
LoRA $\alpha$ 16 16 16 16 16 16
LoRA Dropout 0.1 0.1 0.1 0.1 0.1 0.1
SuperGLUE BoolQ CB RTES MultiRC WiC COPA
Batch size 10 10 10 10 10 10
Epoch 10 20 10 10 10 10
Learning rate 1e-5 5e-5 1e-5 2e-5 1e-5 3e-5
Max sequence length 512 512 512 512 512 256
Gradient accumulation steps 2 2 2 2 2 2
LoRA $r_{q},r_{v}$ 8 8 8 8 8 8
LoRA $\alpha$ 16 16 16 16 16 16
LoRA Dropout 0.1 0.1 0.1 0.1 0.1 0.1
Table 6: The hyperparameters settings for GLUE and SuperGLUE tasks for Llama3.2-3B. For the COPA dataset we implement LlamaForMultipleChoice by ourselves.
@@
Tasks P R F1 Accuracy CON CCON Tasks P R F1 Accuracy CON CCON
MRPC 89.62 87.83 88.59 90.34 92.89 86.79 BoolQ 79.09 80.78 79.71 83.05 86.79 76.37
($\pm$1.39) ($\pm$1.15) ($\pm$1.02) ($\pm$ 0.89) ($\pm$ 1.18) ($\pm$ 0.91) ($\pm$16.87) ($\pm$10.82) ($\pm$14.54) ($\pm$ 7.35) ($\pm$ 12.21) ($\pm$ 10.76)
QNLI 93.98 93.98 93.97 94.00 95.64 91.80 CB 89.43 90.81 89.61 94.64 91.50 89.84
($\pm$0.36) ($\pm$0.36) ($\pm$0.36) ($\pm$ 0.38) ($\pm$ 0.52) ($\pm$ 0.46) ($\pm$ 6.15) ($\pm$ 6.69) ($\pm$ 6.11) ($\pm$ 3.95) ($\pm$ 3.21) ($\pm$ 3.40)
QQP 91.22 91.73 91.46 92.00 95.57 89.79 RTES 71.71 76.49 72.99 72.89 70.05 64.58
($\pm$0.07) ($\pm$0.17) ($\pm$0.07) ($\pm$ 0.06) ($\pm$ 0.30) ($\pm$ 0.16) ($\pm$25.46) ($\pm$14.21) ($\pm$21.72) ($\pm$ 16.67) ($\pm$ 18.36) ($\pm$17.43)
SST2 95.48 95.45 95.46 95.59 96.83 94.01 MultiRC 67.62 74.32 70.05 76.46 74.63 63.77
($\pm$0.42) ($\pm$0.43) ($\pm$0.43) ($\pm$ 0.55) ($\pm$ 0.57) ($\pm$ 0.51) ($\pm$26.94) ($\pm$16.79) ($\pm$23.24) ($\pm$13.30) ($\pm$18.94) ($\pm$15.65)
RTEG 85.41 84.74 84.86 85.02 91.09 80.55 WiC 0.7053 0.6857 0.6778 68.57 80.36 58.75
($\pm$1.45) ($\pm$1.54) ($\pm$1.52) ($\pm$ 1.48) ($\pm$ 1.74) ($\pm$ 1.34) ($\pm$2.68) ($\pm$2.77) ($\pm$3.28) ($\pm$ 2.77) ($\pm$ 6.07) ($\pm$ 4.60)
CoLA - - - 65.61 93.95 82.76 COPA 73.03 72.85 72.78 73.00 64.88 55.44
- - - ($\pm$ 1.32) ($\pm$ 0.71) ($\pm$ 0.51) ($\pm$ 16.69) ($\pm$ 16.59) ($\pm$ 16.81) ($\pm$ 16.83) ($\pm$ 17.35) ($\pm$ 11.98)
Table 7: Additional evaluation performance using RoBERTa-large. P - Precision, R - Recall, F1, Accuracy, CON - consistency, CCON - correct consistency. CoLA employs Matthew’s correlation coefficient as the accuracy.
GLUE SuperGLUE
Tasks P R F1 Accuracy CON CCON Tasks P R F1 Accuracy CON CCON
MRPC 81.98 80.31 81.03 84.02 90.34 79.19 BoolQ 70.78 69.29 69.74 72.49 81.10 63.09
($\pm$ 0.57) ($\pm$1.01) ($\pm$0.79) ($\pm$ 0.56) ($\pm$1.66) ($\pm$0.70) ($\pm$1.77) ($\pm$2.07) ($\pm$2.07) ($\pm$ 1.62) ($\pm$1.67) ($\pm$1.58)
QNLI 94.22 94.22 94.21 94.20 96.85 92.63 CB 56.33 56.23 55.32 73.92 89.82 70.35
($\pm$0.15) ($\pm$0.15) ($\pm$0.14) ($\pm$ 0.14) ($\pm$0.24) ($\pm$0.18) ($\pm$10.07) ($\pm$3.77) ($\pm$5.78) ($\pm$ 2.03) ($\pm$2.92) ($\pm$2.55)
QQP 88.36 89.08 88.69 89.37 96.63 87.69 RTES 68.83 68.82 68.64 68.66 71.69 54.51
($\pm$0.07) ($\pm$0.21) ($\pm$0.12) ($\pm$ 0.09) ($\pm$0.20) ($\pm$0.09) ($\pm$2.84) ($\pm$2.83) ($\pm$2.78) ($\pm$ 2.77) ($\pm$3.82) ($\pm$3.38)
SST2 96.78 96.79 96.78 96.78 98.32 95.95 MultiRC 79.72 80.02 79.83 80.14 86.73 73.50
($\pm$0.27) ($\pm$0.27) ($\pm$0.27) ($\pm$0.27) ($\pm$0.63) ($\pm$0.44) ($\pm$0.57) ($\pm$0.55) ($\pm$0.57) ($\pm$0.57) ($\pm$1.06) ($\pm$0.84)
RTEG 85.92 85.91 85.88 85.92 88.80 80.32 WiC 68.99 68.71 68.60 68.71 85.86 61.64
($\pm$3.36) ($\pm$3.34) ($\pm$3.35) ($\pm$ 3.34) ($\pm$3.05) ($\pm$3.40) ($\pm$1.64) ($\pm$1.58) ($\pm$1.58) ($\pm$ 1.58) ($\pm$1.79) ($\pm$1.08)
CoLA - - - 61.33 94.86 81.36 COPA 84.08 84.05 84.03 84.20 86.8 77.6
- - - ($\pm$ 0.88) ($\pm$0.56) ($\pm$ 0.33) ($\pm$5.25) ($\pm$5.28) ($\pm$5.24) ($\pm$5.17) ($\pm$ 3.99) ($\pm$ 4.83)
Table 8: Additional evaluation performance using Llama3.2-3B. P - Precision, R - Recall, F1, Accuracy, CON - consistency, CCON - correct consistency. CoLA employs Matthew’s correlation coefficient as the accuracy.